# Refining Neural Activation Patterns for Layer-Level Concept Discovery in Neural Network-Based Receivers

### v1

@@ D.2 Model input
D.3 Model architecture
The architecture of the DeepRx model, based on the work by Honkala et al. (2021) and Korpi et al. (2021), is shown in Table 2. It consists of multiple pre-activation ResNet blocks—an extension of the ResNet architecture by He et al. (2016)—and several convolutional layers. The pre-activation ResNet block, illustrated in Figure 10, is composed of two consecutive groups, each consisting of a batch normalization layer, a Rectified Linear Unit (ReLU) layer, and a depthwise separable convolution layer (Chollet (2017)). In each depthwise separable convolution block, the depth multiplier value is set to 1, meaning the number of output channels in the depthwise convolution step is the same as the input. The data flow direction is depicted in Figure 11. Certain architectural details have been omitted as they are not essential to the core contributions of our work.
@@ D.3 Model architecture
LLR Output $\mathbf{L}$, see Section D.4 Output $(14,192,8)$
Figure 10: DeepRx’s pre-activation ResNet block (adapted from Honkala et al. (2021)).
@@
D.4 Model Output and Loss Function
Since DeepRx (Honkala et al. (2021)) is designed to support multiple Quadrature Amplitude Modulation (QAM) schemes in 5G, including QPSK, 16-QAM, 64-QAM, and 256-QAM, which are hierarchically related (see Figure 13).
Figure 12: Modulation schemes used in 5G NR, including BPSK, Pi/2 BPSK, QPSK, 16-QAM, 64-QAM, and 256-QAM (redrawn from Zaidi et al. (2018)).
@@ D.4 Model Output and Loss Function
Figure 14: The relationship between 16-QAM and QPSK constellation points (redrawn from Honkala et al. (2021)). When QPSK is used, only 2 bits $b_{1}$ and $b_{2}$ are used to define the complex quadrant, but if 16-QAM is used, besides the first two bits, $b_{3}$ and $b_{4}$ are also used to define the quadrant within the quadrant.
Figure 15: Masking bits based on the modulation, including both QPSK and 16QAM, where masking is selected based on the modulation. The output is masked accordingly so that unnecessary parts of the output array are not considered (redrawn from Honkala et al. (2021)).
@@
$\text{CE}(\bm{\theta})=-\frac{1}{\#DB}\sum_{(i,j)\in\mathcal{D}}\sum_{l=0}^{B-1}\left(b_{ijl}\log(\hat{b}_{ijl})+(1-b_{ijl})\log(1-\hat{b}_{ijl})\right),$ (9)
where #$\#\mathcal{D}$ is the number of resource elements carrying data; $B$ is the number of bits in the constellation used ($B=4$ for 16-QAM); $b_{ijl}$ is the label bit; and $\hat{b}_{ijl}$ is the probability that bit $b_{ijl}$ equals one, calculated using the sigmoid function:
@@
D.5 Model’s Predictive Performance
@@
The DeepRx model demonstrates competitive performance in terms of uncoded Bit Error Rate (BER) across varying Signal-to-Interference-plus-Noise-Ratio (SINR) levels, as seen in Figure 13. It closely matches the baseline performance with a known channel, particularly at low SINR values. Notably, it outperforms the baseline with one pilot in higher SINR scenarios, which aligns with the findings reported in Honkala et al. (2021).
Appendix E Supplementary Analysis of Dataset Size Impact on Visual Object Recognition Using ResNet50 with Pretrained Weights and ImageNet Data
This appendix explores the impact of dataset size on visual object recognition using the ResNet50 architecture with pretrained weights and activations from the Conv4 and Conv5 stages of the network, as described in Appendix C. Figures 17 and 17 illustrate how varying the dataset size influences the number of clusters and the mean persistence score across different layers and methodologies.
Appendix F Supplementary Analysis of Dataset Size Impact on Radio Receiver Using DeepRx with Simulated Data
This appendix explores the impact of dataset size on radio receiver using the DeepRx architecture and activations from various stages of the network, as described in Appendix D. Figures 19 and 19 illustrate how varying the dataset size influences the number of clusters and the mean persistence score across different layers and methodologies.
Appendix G Experiments on Delay Spread and UE Speed Using the NUCPI Methodology
@@
In this appendix, we present our replication of the experiments conducted by Tuononen et al. (2025), employing delay spread (see Appendix G.2) and user equipment (UE) speed (see Appendix G.3) as target parameters. This approach diverges from the original study, which utilized Signal-to-Noise Ratio (SNR) as the channel parameter of interest. Notably, Tuononen et al. (2025) suggested that their method could be applicable to other channel parameters, which we explore here.
G.1 NUCPI Methodology and Its Potential
@@
We refer to the methodology proposed in Tuononen et al. (2025) as Neural Unit Channel Parameter Identification (NUCPI), wherein an explainer model is trained to predict a given channel parameter based on intermediate activations from a performer model—in our case, the DeepRx model (see Appendix D). We interpret the predictive performance of the explainer model—measured using Mean Squared Error (MSE), as suggested by Tuononen et al. (2025)—as an indicator of how much information specific internal components (e.g., layers or channels) of the performer contain about the given channel parameter. This enables insights into the model’s internal behavior.
@@
These insights could potentially be used to identify redundant or over-specialized components, guide model pruning Han et al. (2016), or adapt architectures to better handle specific channel conditions Cortes et al. (2017). While these applications are not experimentally explored in this work, the interpretability provided by NUCPI opens promising directions for future research in model compression and adaptive architecture design.
G.2 NUCPI Analysis of Delay Spread
We performed NUCPI analysis of delay spread using both layerwise and channelwise (for layer B1-PRE) perspectives. As shown in the layerwise (Figure 20) and channelwise (Figure 21) test losses of the explainer models, predictive performance is consistently poor across all units. Upon closer examination, it is evident that the explainer models are overfitting: while training loss decreases steadily, evaluation loss plateaus or increases, indicating limited generalization to unseen data (Goodfellow et al. (2016)). This is illustrated in training curves for both the layerwise (B1-PRE in Figure 23; OUT-POST in Figure 23) and channelwise (channels 7 and 12 on B1-PRE, in Figures 25 and 25, respectively) explainer models.
From a physical layer perspective, this is plausible given our simulation parameters (see Table 1). The bandwidth is $B=192\times 30\text{ kHz}=5.76\text{ MHz}$, yielding a time resolution of $\Delta\tau=1/B\approx 174\text{ ns}$, and a quantization uncertainty of approximately $\sigma\approx\Delta\tau/\sqrt{12}\approx\pm 50\text{ ns}$ (standard deviation of a uniform distribution (Earl and Nicholson, 2021, Appendix 15)). As a result, multipath components closer than 174 ns apart (e.g., a 10 ns path) are indistinguishable, and even larger separations are affected by quantization noise. This limitation likely contributes to the lack of delay spread information in DeepRx’s activations. Interestingly, DeepRx still performs well under delay spread conditions, suggesting it may compensate for this effect without needing to explicitly quantify the delay spread (e.g., as RMS).
G.3 NUCPI Analysis of User Equipment (UE) Speed
We also applied NUCPI analysis to UE speed, evaluating layerwise explainer performance for two layers. As with delay spread (see Appendix G.2), the explainer models exhibit consistent overfitting: training losses decrease, but evaluation losses do not improve, as shown in Figures 27 and 27. Based on this behavior, we conclude that DeepRx’s internal activations likely do not encode UE speed in a directly predictive way.
In traditional systems, UE speed—or Doppler spread more technically—is estimated by aggregating over tens of slots (Yucek et al. (2005); Zhao et al. (2009)). In contrast, DeepRx operates on a single slot without a memory of past slots. Despite this, DeepRx can still handle frequency offset effects caused by user movement (and other factors). A similar explanation as with delay spread may apply: the model may not need to explicitly estimate frequency offset or Doppler spread to effectively compensate for them.

### v2

@@ B.2 Model Input
B.3 Model Architecture
The architecture of the DeepRx model, based on the work by honkala2021deeprx and korpi2021, is shown in Table 8. It consists of multiple pre-activation ResNet blocks—an extension of the ResNet architecture by he2016resnet—and several convolutional layers. The pre-activation ResNet block, illustrated in Figure 16, is composed of two consecutive groups, each consisting of a batch normalization layer, a Rectified Linear Unit (ReLU) layer, and a depthwise separable convolution layer (Chollet2017). In each depthwise separable convolution block, the depth multiplier value is set to 1, meaning the number of output channels in the depthwise convolution step is the same as the input. The data flow direction is depicted in Figure 17. Certain architectural details have been omitted as they are not essential to the core contributions of our work.
@@ B.3 Model Architecture
LLR Output $\mathbf{L}$, see Section B.4 Output $(14,192,8)$
Figure 16: DeepRx’s pre-activation ResNet block (adapted from honkala2021deeprx).
@@
B.4 Model Output and Loss Function
Since DeepRx (honkala2021deeprx) is designed to support multiple Quadrature Amplitude Modulation (QAM) schemes in 5G, including QPSK, 16-QAM, 64-QAM, and 256-QAM, which are hierarchically related (see Figure 18).
Figure 18: Modulation schemes used in 5G NR, including BPSK, Pi/2 BPSK, QPSK, 16-QAM, 64-QAM, and 256-QAM (redrawn from Zaidi2018).
@@ B.4 Model Output and Loss Function
Figure 20: The relationship between 16-QAM and QPSK constellation points (redrawn from honkala2021deeprx). When QPSK is used, only 2 bits $b_{1}$ and $b_{2}$ are used to define the complex quadrant, but if 16-QAM is used, besides the first two bits, $b_{3}$ and $b_{4}$ are also used to define the quadrant within the quadrant.
Figure 21: Masking bits based on the modulation, including both QPSK and 16QAM, where masking is selected based on the modulation. The output is masked accordingly so that unnecessary parts of the output array are not considered (redrawn from honkala2021deeprx).
@@
$\text{CE}(\bm{\theta})=$ $-\frac{1}{\#DB}\sum_{(i,j)\in\mathcal{D}}\sum_{l=0}^{B-1}$
$\left(b_{ijl}\log(\hat{b}_{ijl})+(1-b_{ijl})\log(1-\hat{b}_{ijl})\right),$ (32)
where $\mathcal{D}$ is the number of resource elements carrying data; $B$ is the number of bits in the constellation used ($B=4$ for 16-QAM); $b_{ijl}$ is the label bit; and $\hat{b}_{ijl}$ is the probability that bit $b_{ijl}$ equals one, calculated using the sigmoid function:
@@
B.5 Model’s Predictive Performance
@@
The DeepRx model demonstrates competitive performance in terms of uncoded Bit Error Rate (BER) across varying Signal-to-Interference-plus-Noise-Ratio (SINR) levels, as seen in Figure 19. It closely matches the baseline performance with a known channel, particularly at low SINR values. Notably, it outperforms the baseline with one pilot in higher SINR scenarios, which aligns with the findings reported in honkala2021deeprx.
Appendix C Neural Unit Channel Parameter Identification (NUCPI) Methodology for Mechanistic Interpretability
C.1 Introduction
@@
We present our replication of the experiments conducted by tuononen2025interpreting, employing delay spread (see Appendix C.3.1) and user equipment (UE) speed (see Appendix C.3.2) as target parameters. This approach diverges from the original study, which utilized Signal-to-Noise Ratio (SNR) as the channel parameter of interest. Notably, tuononen2025interpreting suggested that their method could be applicable to other channel parameters, which we explore here.
C.2 Methodology
@@
We refer to the methodology proposed in tuononen2025interpreting as Neural Unit Channel Parameter Identification (NUCPI), wherein an explainer model is trained to predict a given channel parameter based on intermediate activations from a performer model—in our case, the DeepRx model (see Supplemental B). We interpret the predictive performance of the explainer model—measured using Mean Squared Error (MSE), as suggested by tuononen2025interpreting—as an indicator of how much information specific internal components (e.g., layers or channels) of the performer contain about the given channel parameter. This enables insights into the model’s internal behavior.
@@
These insights could potentially be used to identify redundant or over-specialized components, guide model pruning han2016deepcompression, or adapt architectures to better handle specific channel conditions cortes2017adanet. While these applications are not experimentally explored in this work, the interpretability provided by NUCPI opens promising directions for future research in model compression and adaptive architecture design.
@@
C.3 Experiments on Radio Receiver
C.3.1 NUCPI Analysis of Delay Spread
We performed NUCPI analysis of delay spread using both layerwise and channelwise (for layer B1-PRE) perspectives. As shown in the layerwise (Figure 22) and channelwise (Figure 23) test losses of the explainer models, predictive performance is consistently poor across all units. Upon closer examination, it is evident that the explainer models are overfitting: while training loss decreases steadily, evaluation loss plateaus or increases, indicating limited generalization to unseen data goodfellow2016deep. This is illustrated in training curves for both the layerwise (B1-PRE in Figure 24; OUT-POST in Figure 25) and channelwise (channels 7 and 12 on B1-PRE, in Figures 26 and 27, respectively) explainer models.
From a physical layer perspective, this is plausible given our simulation parameters (see Table 1). The bandwidth is $B=192\times 30\text{ kHz}=5.76\text{ MHz}$, yielding a time resolution of $\Delta\tau=1/B\approx 174\text{ ns}$, and a quantization uncertainty of approximately $\sigma\approx\Delta\tau/\sqrt{12}\approx\pm 50\text{ ns}$ (standard deviation of a uniform distribution (earl2021_appendix15, Appendix 15)). As a result, multipath components closer than 174 ns apart (e.g., a 10 ns path) are indistinguishable, and even larger separations are affected by quantization noise. This limitation likely contributes to the lack of delay spread information in DeepRx’s activations. Interestingly, DeepRx still performs well under delay spread conditions, suggesting it may compensate for this effect without needing to explicitly quantify the delay spread (e.g., as RMS).